Chapter 6, Section 4-D, Exercise 184 Use a t-distribution to answer this question. Assume the samples are random samples from distributions that are reasonably normally distributed, and that a t-statistic will be used for inference about the difference in sample means. State the degrees of freedom used. Find the endpoints of the-distribution with 2.5% beyond them in each tail if the samples have sizes ni-17 and n2-27. Enter the exact answer for the degrees of freedom and round your answer for the endpoints to two decimal places. degrees of freedom = endpoints+
Answer #1

Solution :

Given that,

n1 = 17 , n2 = 27

Degrees of freedom = n1 + n2 - 2 = 17 + 27 - 2 = 42

t\alpha /2,df = +/-2.02